OpenStudy (yeval76):

****FAN AND MEDAL**** Which of the following expressions represents a function? x = 1 {(3, 2), (3, –2), (4, 5), (4, –5)} y = 4x − 1 4x2 + y2 = 16

OpenStudy (yeval76):

@ganeshie8

OpenStudy (yeval76):

@hartnn

OpenStudy (yeval76):

Help Please!!!

OpenStudy (yeval76):

@ganeshie8 @hartnn pLEASE hELP

OpenStudy (calculusxy):

A function has an input and only one output.

OpenStudy (yeval76):

Ok so we can take out the second one right?

OpenStudy (calculusxy):

Yes

OpenStudy (calculusxy):

One type of function would be a linear function: y=mx + b

OpenStudy (calculusxy):

Don't get fooled by the plus sign because if b is a negative integer, then it would be like y=mx + (-b) or y=mx - b. It means the same thing.

OpenStudy (calculusxy):

Do you see anything like that in the choices?

OpenStudy (calculusxy):

@yeval76